experiment indicates that for long wave lengths that the speed c is effectively independent of amplitude and surface, suppose c=kg^xλp^z where k is a dimensionless constant, show by dimensional analysis that c is √kgλ?
1
Expert's answer
2021-08-02T14:33:06-0400

Gives


"c=kg^x\lambda p^z\\ L.H.S. demension \\c=[M^0L^1T^{-1}]\\R.H.S dimensions \\=[M^0L^1T^{-2}]^{x}[M^0L^1T^0][ML^{-1}T^{-2}] ^z"

"[M^0L^1T^{-2}]=[M]^{z}[L]^{x+1-z}[T]^{-2x-2z}"

Comparison x y ,z pawor

"1=x-1-z"

"z=0"

"-1=-2x-2z"

"x=\frac{1}{2}"

Put x ,y value

"c=\sqrt{kg\lambda}"

Hence proved